OpenStudy (anonymous):

Determine the value of A so that the line whose equation is Ax+y-5=0 is perpendicular to the line containing the points (1,-4) and(-5,4). A=?

OpenStudy (gorv):

slope of perpendicular line m=(4-(-4))/(-5-1)=8/-6=-4/3

OpenStudy (gorv):

so slope of other line must be =-1/(-4/3)=3/4 property of perpendicular lines

OpenStudy (gorv):

ax+y-5=0 comverting in standard form y=m1*x+c y-5=-ax y=-ax+5 on comparing m1=-a -a=3/4 a=-3/4

OpenStudy (anonymous):

@gorv how did you get y-5=-ax from y=ml*x+c?

OpenStudy (anonymous):

@gorv nvm i got it but our teacher never taught us those formulas : y=ml*x+c or ml=-a

OpenStudy (anonymous):

I see why thats the answer though... still frustrating because I would of been driving in circles if I hadn't known the formula y=ml*x+c

OpenStudy (gorv):

you heard of y=mx+c

OpenStudy (gorv):

??

OpenStudy (gorv):

reply plzzz

OpenStudy (gorv):

as i m getting no rply ..so i m leaving it here ,,,